# Observing hidden neuronal states in experiments

### v2

@@ Abstract
In this article, we revisit voltage-clamp electrophysiology, a feedback-control method that has successfully been applied to study neuronal (non)linear current-voltage relationships. Yet its full scope has not been explored to date. We reinterpret the voltage-clamp protocol under slow variation of the feedback reference signal using the language of multiple-timescale dynamical systems theory. Gauged under this new epistemic step, we unveil non-observable (hidden) states. More specifically, the experimental measurements from voltage clamp can be put into correspondence with predictions of both stable and unstable states from dynamical systems theory. We demonstrate a general protocol for constructing systematically experimental steady-state bifurcation diagrams for electrophysiologically active cells. We perform our experiments on entorhinal cortex neurons, both excitatory (pyramidal neurons) and inhibitiory (interneurons). Moreover, employing a current-clamp open-loop electrophysiology protocol with slow variation of an applied electrical current drives the neuron to dynamically transition between its observable rest states and its observable spiking states, hence revealing experimental dynamic bifurcations. Overlaying the two aforementioned protocols produces a map of all these neuronal states and effectively leads to an experimental validation of the classical slow-fast dissection method introduced by J. Rinzel in the 1980s and routinely applied ever since in order to analyse slow-fast neuronal models. Our approach opens doors to observing further complex hidden states with more advanced control strategies, allowing to control real cells beyond pharmacological manipulations.
Introduction.–
@@
When characterising the dynamics of nonlinear systems, one fundamental criterion for a model is if its stable states such as stationary solutions or periodic orbits match experimental observations. The ability to validate models is, thus, greatly expanded by experimental tools with the capacity to unveil non-observable (sensitive or dynamically unstable) states that are otherwise inaccessible to standard measurements. The combination of observable and non-observable states gives access to an experimental equivalent of parameter-dependent families of stable and unstable states in a model, which are usually referred to as a bifurcation diagram.
@@
This article applies a new experimental technique of using feedback control to find unstable states to electrophysiology experiments on neuronal cells. Our aim is to support systematic validation of neuron models by comparing bifurcation diagrams and observing their between-cell variability. We focus on unstable parts of input-dependent families of steady-state solutions obtained by feedback-controlled experiments and compare them with indirect evidence from standard measurements from open-loop experiments. This extends recent work of Ori et al. [1, 2] constructing phase diagrams from neuronal data, and complements other approaches such as using data to verify the bifurcation structure of neuronal models [3, 4], model-based data analysis [5] or parameter estimation from data [6].
@@
Our technique is simplification of the so-called control-based continuation (CBC) method, an approach which has been recently demonstrated in mechanical experiments [7], vibrations and buckling experiments [8], pedestrian flow experiments [9], atomic-force microscopy [10], cylindrical pipe flow simulations [11], and feasibility studies for synthetic gene networks [12, 13]. Indeed, CBC is a procedure that combines feedback control [14] and pseudo-arclength continuation [15] in a model-free environment, that is, only reliant upon noisy experimental data in a closed-loop control setup. The objective is to compute experimental bifurcation diagrams, that is, both families of stable and unstable states (either stationary or periodic) together with bifurcation points joining such families. Importantly, the control is noninvasive in the sense that it vanishes once an equilibrium or periodic of the uncontrolled system has been reached; the control is decreased iteratively using a Newton’s method. In our case, the approach is simpler and does not require the use of Newton iterations since we consider a slow ramp on the control target. Hence we exploit de presence of two timescales in the experimental procedure in order to obtain directly an approximation of the steady-state bifurcation curve, and we explain why such a bifurcation curve is obtained at low cost; see below.
@@
In our electrophysiology experiments on entorhinal cortex neurons we treat the cell as an electric circuit, and apply a voltage clamp (VC) [16, 17], followed by current clamp (CC). In the VC setup the electrode acts as a voltage source at the neural membrane, fixing the potential across the neural membrane, measuring the current, while the CC setup adds a fixed external current, measuring the resulting membrane potential; see Fig. 1(a) and the Supplementary Information (SI) for further details. In-vivo neurons are subject to current signals that drive spiking (oscillatory) or rest (steady) states of the neural membrane potential [18]. This makes the CC setup the open-loop part of the protocol. In contrast, the VC part of our experiment the closed-loop feedback-controlled part of the protocol because the voltage source regulates the external current to maintain the hold voltage. VC has been applied successfully to study neuronal nonlinear current-voltage relationships, so-called N-shaped $I$-$V$ characteristics, which cause enhanced neuronal excitability and influence the regenerative activation of certain ionic currents (e.g., sodium) [19, 20, 21, 22, 23] across the neural membrane, into and out of the cell. We show that the VC protocol with a slowly varied reference voltage signal gives access to stable and unstable neuronal steady states of the neuron, which was hinted at in [24, 25]. In contrast, the open loop CC protocol with a slowly varying applied current always follows stable (observable) states, driving the neuron to dynamically transition between its observable rest states and its observable spiking states.
@@
We interpret these combined experimental protocols (VC and CC) using multiple-timescale dynamics, in particular, the dissection method [26], which reveals the dynamic bifurcations in the experiment; see Fig. 1(b).
@@
Figure 1: (a) Experimental setup with brain slice (A), patch pipette (B), reference electrode ((Ag-AgCl pellet) connected to ground (C), amplifier (E, Multiclamp 700B) with CV-7B headstage (D), AD-converter (F, National Instruments NI USB-6343) and standard PC computer (G). (b) VC and CC protocol runs for cell $5$: $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$-curve for VC run (thin bright blue: unfiltered data with sampling time step $5\times 10^{-5}$ s, blue/red/brown: median of $I_{\mathrm{vc}}$ over moving windows of size $\Delta w=4\times 10^{3}$ steps equalling $0.2$s) and $(I_{\mathrm{h}}(t),V_{\mathrm{cc}}(t))$-curve for CC run (orange, mean of $I_{\mathrm{h}}$ over moving windows. (c) $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$-curves for VC protocol of all $5$ cells on waterfall $V_{\mathrm{h}}$-axis (same color coding as (b)).
@@
In this slow-fast framework, the states traced by the VC protocol with slow variation correspond to the steady-state experimental bifurcation diagram of the so-called fast subsystem of the mathematical model describing the protocol [26], that is, the model system with constant input current. Hence, the N-shaped I-V relation of a neuron should be seen as a S-shaped V-I bifurcation diagram; see again Fig. 1(b).
@@
Results.–
@@
Figure 1(b) shows the time profiles $(I_{\mathrm{h}}(t),V_{\mathrm{cc}}(t))$ of the CC protocol run (orange, thin) and $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$ of the VC protocol run (bright blue, thin) for cell $5$ overlaid in the $(I,V)$-plane. After smoothing the VC time profile is the S-shaped curve $(I_{\mathrm{vc,sm}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ (blue/brown/red, thick). It equals the $(I,V)$-characteristic of the stationary neuronal states of the CC protocol, including dynamically unstable states (brown and red). The transition to stable spiking states is compatible with type-I excitability, however we do not have sufficient data to distinguish the different types of excitability; see SI, Fig. S2. Their dynamical instability is inferred in two ways: (i) by the negative slope of the $(I_{\mathrm{vc,sm}},V_{\mathrm{h}})$-curve (brown) after smoothing over a moving window with larger size $\Delta w=1.5\times 10^{4}$ steps ($=0.75$s), or (ii) by the presence of oscillations (neuronal firing) in the CC run at $I_{\mathrm{h}}$ equalling the $I_{\mathrm{vc,sm}}$ (red). The stability boundaries of the stationary states are labelled as bifurcations in Fig. 1(b). The change of stability near the disappearance of stable spiking states at $I_{\mathrm{db}}$ is a Hopf bifurcation. The fold points of the $(I_{\mathrm{vc,sm}},V_{\mathrm{h}})$-curve are saddle-node (fold) bifurcations. Figure 1(c) shows the stationary-state curves with their inferred dynamical stability for all $5$ cells (see SI for CC run time profiles used to partially infer stability of cells 11114444). The cells are vertically ordered and numbered according to depth of the S-shape, determining if they fall into the class of type-I or type-II neurons. Figure 1(c) demonstrates wide variability in steady-state curve shape among cells of nominally same function.
@@
Figure 2: (a) Experimental bifurcation diagrams for an interneuron from the entorhinal cortex. (b) Experimental bifurcation diagrams for a type-II PY neuron from the same region. Protocols identical to Fig. 1(b).
What is more, the method described in this work is versatile and rests upon only standard electrophysiological procotols. Hence, it can be applied to any neuron type. Fig. 2(a) below illustrates this point by showing the results of the VC and CC protocols applied to an inhibitory neuron from the entorhinal cortex while panel (b) shows that the same protocols can be applied to a type-II neuron. The steady-state curve is seemingly more complicated than for PY neurons’ recordings shown on other figures in this supplemental material, as well as, in the main manuscript, however it is compatible with a type-I excitable neuron. Noteworthy, there are examples of such scenarios in the literature, in particular with low-threshold spiking neurons [27]. We also have recordings (not shown) from the PY neurons of the dentate gyrus showing a similar outcome once the result of the VC protocol is superimposed onto the result of the CC protocol.
Analysis.–
@@
describing the current balance across the neuron’s membrane. The membrane potential is $V$, $I_{j}(x_{j},V)_{j=1,\ldots,N}$ are the currents across different voltage-gated ion channel types and $I_{\mathrm{ext}}$ is the external current. Each channel type $j$ has a set of associated gating variables $x_{j}$, which are possibly vectors of length $n_{j}$ if the channel gate has both activation and inactivation states, with steady-state gating functions $x_{\infty,j}(V)$ (also of size $n_{j}$) and relaxation times $\tau_{j}(V)$ (a diagonal $n_{j}\times n_{j}$ matrix). The observed dynamic effects such as oscillations (firing/spiking) and negative-slope $(I,V)$ characteristics are determined by these channel coefficients $I_{j}$, $x_{\infty,j}$ and $\tau_{j}$ that are traditionally obtained by parameter fitting from VC experiments, a difficult and ill-posed problem, as gating variables $x_{j}$ are not directly measured [28].
@@
The VC and CC protocols use different mechanisms for generating $I_{\mathrm{ext}}(t)$. The VC protocol is a closed loop where a voltage source regulates $I_{\mathrm{ext}}$ with high gain $g_{\mathrm{c}}$ to achieve the slowly varying hold voltage $V_{\mathrm{h}}$ at the voltage source for (1), measuring $I_{\mathrm{vc}}$:
$\begin{split}I_{\mathrm{ext}}&\approx I_{\mathrm{vc}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V),\\ \dot{V}_{\mathrm{h}}&=\varepsilon\Delta_{V}\end{split}$ (2)
@@
which turns (1), (2) into a multiple-timescale dynamical system with $1+\sum n_{j}$ fast state variables $(V,x_{j})$ (where $\sum n_{j}$ is the overall number of gating variables) and one slow state variable $V_{\mathrm{h}}$, corresponding to the feedback reference signal [29]. The speed at which $V_{\mathrm{h}}$ varies is $\varepsilon\Delta_{V}$ with $\Delta_{V}(t)=0.183$ mV///ms, where we extract the dimensionless small factor $\varepsilon=10^{-2}$.
@@
In contrast, the CC protocol holds $I_{\mathrm{ext}}$, measuring the generated voltage $V_{\mathrm{cc}}$, thus, corresponding to an open-loop system, permitting e.g. the spiking seen in Fig. 1(b):
$\begin{split}I_{\mathrm{ext}}&\approx I_{\mathrm{h}},\\ V_{\mathrm{cc}}&\approx V,\\ \dot{I}_{\mathrm{h}}&=\varepsilon\Delta_{I}.\end{split}$ (3)
@@
The applied hold current $I_{\mathrm{h}}$ is varied slowly at speed $\varepsilon\Delta_{I}$ with $\varepsilon=10^{-2}$ and $\Delta_{I}=0.66\ldots 0.75$ pA///ms. System (1), (3) is also a slow-fast system with $1+N$ fast variables $(V,x_{j})$ and $1$ slow variable $I_{\mathrm{h}}$. The gain $g_{\mathrm{c}}$ in (2) is limited by the imperfect conductance across the non-zero spatial extent of the membrane. Even though (1) is for the potential $V$ across the entire membrane and only $V_{\mathrm{h}}$ at the clamp is measured, we approximate $V_{\mathrm{h}},V_{\mathrm{cc}}\approx V$ for, the membrane potential, and $I_{\mathrm{h}},I_{\mathrm{vc}}\approx I_{\mathrm{ext}}$ for the external current in (1).
@@
Following a classical multiple-timescale approach, we consider the $\varepsilon=0$ limit of system (1), (2), which corresponds to its $(1+\sum n_{j})$-dimensional fast subsystem (1), with $I_{\mathrm{ext}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$, where $V_{\mathrm{h}}$ is now treated as a parameter. For fixed $V_{\mathrm{h}}$ and voltages in the range $-80\ldots 30$ mV of interest, system (1), (2) has only stable steady states (no limit cycles, that is, no neuronal spikes); see SI for details. For a fixed hold voltage $V_{\mathrm{h}}$, the steady states of (1), (2) with $I_{\mathrm{ext}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$ satisfy the algebraic equations for $(V,x_{j})_{j=1,\ldots,N}$
$\begin{split}\textstyle{\sum_{j}}I_{j}(x_{j},V)&=g_{\mathrm{c}}(V_{\mathrm{h}}-V),\\ x_{j}&=x_{\infty,j}(V),\end{split}$ (4)
@@
where $I_{\mathrm{eq}}(V)=\sum_{j}I_{j}(x_{\infty,j},V)$ is the equilibrium current for fixed membrane potential $V$. The solutions of (4) form a (1D) steady-state curve of system (1), (2), which is normally hyperbolic (transversally attracting) for $\varepsilon=0$. For $\varepsilon\neq 0$ the increase of $V_{\mathrm{h}}$ with speed $\varepsilon\Delta_{V}$ introduces a slow variation of all states $(V,x_{j})$. Hence, $V$ and the feedback current $I_{\mathrm{vc}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$ (as measured), are not at their steady-state values given by (4), but they are changing dynamically. This results in a difference between the measured curve $(I_{\mathrm{vc}},V_{\mathrm{h}})$ in Fig. 1(b) and the $(I,V)$-values of the desired steady-state curve (4).

### v3

@@ Abstract
In this article we demonstrate a general protocol for constructing systematically experimental steady-state bifurcation diagrams for electrophysiologically active cells. We perform our experiments on entorhinal cortex neurons, both excitatory (pyramidal neurons) and inhibitiory (interneurons). A slowly ramped voltage-clamp electrophysiology protocol serves as closed-loop feedback controlled experiment for the subsequent current-clamp open-loop protocol on the same cell. In this way, the voltage-clamped experiment determines dynamically stable and unstable (hidden) steady states of the current-clamp experiment. The transitions between observable steady states and observable spiking states in the current-clamp experiment provide partial evidence for stability and bifurcations of the steady states. This technique for completing steady-state bifurcation diagrams in a model-independent way expands support for model validation to otherwise inaccessible regions of the phase space. Overlaying the voltage-clamp and current-clamp protocols leads to an experimental validation of the classical slow-fast dissection method introduced by J. Rinzel in the 1980s and routinely applied ever since in order to analyse slow-fast neuronal models. Our approach opens doors to observing further complex hidden states with more advanced control strategies, allowing to control real cells beyond pharmacological manipulations.
1Laboratory of Molecular Mechanisms of Neural Interactions,
Sechenov Institute of Evolutionary Physiology and Biochemistry of RAS, Saint Petersburg, Russia
2Institute for Theoretical Physics, University of Bremen, Bremen, Germany
3MCEN Team, Basque Center for Applied Mathematics (BCAM), Bilbao, Spain
4MathNeuro Project-Team, Inria Branch of the University of Montpellier, Montpellier, France
5College of Engineering, Mathematics and Physical Sciences, University of Exeter, Exeter, United Kingdom
6Ikerbasque, the Basque Foundation for Science, Bilbao, Spain
* mathieu.desroches@inria.fr; Jan.Sieber@exeter.ac.uk; srodrigues@bcamath.org
1 Introduction
@@
When characterising the dynamics of nonlinear systems, one fundamental criterion for a model is if its stable states such as stationary solutions or periodic orbits match experimental observations. The ability to fit and validate models is, thus, greatly expanded by experimental tools with the capacity to unveil non-observable (sensitive or dynamically unstable) states that are otherwise inaccessible to standard measurements. The combination of observable and non-observable states gives access to an experimental equivalent of parameter-dependent families of stable and unstable states in a model, which are usually referred to as a bifurcation diagram.
@@
This article applies a new experimental technique of using feedback control to find unstable states to electrophysiology experiments on neuronal cells. Our aim is to support systematic validation of neuron models by comparing bifurcation diagrams and observing their between-cell variability. We focus on unstable parts of input-dependent families of steady-state solutions obtained by feedback-controlled experiments and compare them with indirect evidence from standard measurements from open-loop experiments. This extends recent work of Ori et al. [ori2018, ori2020] constructing phase diagrams from neuronal data, and complements other approaches such as using data to verify the bifurcation structure of neuronal models [levenstein2019, hesse2022], model-based data analysis [sip2023] or parameter estimation from data [ladenbauer2019].
@@
Our technique is a simplification of the so-called control-based continuation (CBC) method, an approach which has been recently demonstrated in mechanical experiments [sieber2008a], vibrations and buckling experiments [RENSON2019449], pedestrian flow experiments [panagiotopoulos2022control], atomic-force microscopy [bottcher2025exposing], cylindrical pipe flow simulations [willis_duguet_omelchenko_wolfrum_2017], and feasibility studies for synthetic gene networks [de2022control, blyth2023numerical]. Indeed, CBC is a procedure that combines feedback control [phillips1999] and pseudo-arclength continuation [allgower2003] in a model-free environment, that is, only reliant upon noisy experimental data in a closed-loop control setup. The objective is to compute experimental bifurcation diagrams, that is, both families of stable and unstable states (either stationary or periodic) together with bifurcation points joining such families. Importantly, the control is noninvasive in the sense that it vanishes once an equilibrium or periodic of the uncontrolled system has been reached; the control is decreased iteratively using a Newton’s method. In our case, the approach is simpler and does not require the use of Newton iterations since we consider a slow ramp on the control target. Hence we exploit de presence of two timescales in the experimental procedure in order to obtain directly an approximation of the steady-state bifurcation curve, and we explain why such a bifurcation curve is obtained at low cost; see below. Our approach is hence useful for both modelers and experimentalists, as it relies upon standard protocols routinely used in patch-clamp electrophysiology. It can further decipher the excitability class of a given real neuron as well as help fitting a model to data by using both current-clamp and voltage-clamp protocols.
@@
In our electrophysiology experiments on entorhinal cortex neurons we treat the cell as an electric circuit, and apply a voltage clamp (VC) [cole1955, hodgkin1952b], followed by current clamp (CC). In the VC setup the electrode acts as a voltage source at the neural membrane, fixing the potential across the neural membrane, measuring the current, while the CC setup adds a fixed external current, measuring the resulting membrane potential; see the illustration for the experimental setup in Fig 1A and the S1 Text for further details. In-vivo neurons are subject to current signals that drive spiking (oscillatory) or rest (steady) states of the neural membrane potential [schaefer2006]. This makes the CC setup the open-loop part of the protocol. In contrast, the VC part of our experiment is the closed-loop feedback-controlled part of the protocol because the voltage source regulates the external current to maintain the hold voltage. VC has been applied successfully to study neuronal nonlinear current-voltage relationships, so-called N-shaped II-$V$ characteristics, which cause enhanced neuronal excitability and influence the regenerative activation of certain ionic currents (e.g., sodium) [schwindt1981, fishman1969, johnston1980, blatt1988, vervaeke2006] across the neural membrane, into and out of the cell. We show that the VC protocol with a slowly varied reference voltage signal gives access to stable and unstable neuronal steady states of the neuron, which was hinted at in [delnegro1998, qian2014]. In contrast, the open loop CC protocol with a slowly varying applied current always follows stable (observable) states, driving the neuron to dynamically transition between its observable rest states and its observable spiking states.
@@
We interpret these combined experimental protocols (VC and CC) using multiple-timescale dynamics, in particular, the dissection method [rinzel87], which reveals the dynamic bifurcations in the experiment, as demonstrated in the experimental bifurcation diagram in Fig 1B.
@@
Figure 1: A: Sketch of experimental setup with brain slice (a), patch pipette (b), reference electrode (Ag-AgCl pellet) connected to ground (c), amplifier (e, Multiclamp 700B) with CV-7B headstage (d), AD-converter (f, National Instruments NI USB-6343) and standard PC computer (g, https://pixabay.com/vectors/computer-desktop-workstation-office-158675/). B: VC and CC protocol runs for cell 55: $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$-curve for VC run (thin bright blue: unfiltered data with sampling time step $5\times 10^{-5}$ s, blue/red/brown: median of $I_{\mathrm{vc}}$ over moving windows of size $\Delta w=4\times 10^{3}$ steps equalling $0.2$s) and $(I_{\mathrm{h}}(t),V_{\mathrm{cc}}(t))$-curve for CC run (orange, mean of $I_{\mathrm{h}}$ over moving windows. C: $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$-curves for VC protocol of all $5$ cells on waterfall $V_{\mathrm{h}}$-axis (color coding indicates conjectured stability as indicated for (b)).
@@
In this slow-fast framework, the states traced by the VC protocol with slow variation correspond to the steady-state experimental bifurcation diagram of the so-called fast subsystem of the mathematical model describing the protocol [rinzel87], that is, the model system with constant input current. Hence, the N-shaped I-V relation of a neuron should be seen as a S-shaped V-I bifurcation diagram, such as in Fig 1B.
@@
2 Results
The experimental neuronal bifurcation diagram in Fig 1B shows the time profiles $(I_{\mathrm{h}}(t),V_{\mathrm{cc}}(t))$ of the CC protocol run (orange, thin) and $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$ of the VC protocol run (bright blue, thin) for cell $5$ overlaid in the $(I,V)$-plane. For both runs $I_{\mathrm{h}}(t)$ and $V_{\mathrm{h}}(t)$ were slowly increased, respectively (see Materials and Methods). After smoothing, the VC time profile is the S-shaped curve $(I_{\mathrm{vc,sm}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ (blue/brown/red, thick). It equals the $(I,V)$-characteristic of the stationary neuronal states of the CC protocol, including dynamically unstable states (brown and parts of red). The transition to stable spiking states is compatible with (Hodgkin) class-I excitability, however we do not have sufficient data to distinguish the different classes of excitability. See Fig B in S1 Text for how the interspike intervals depend on $I_{\mathrm{h}}$.
The dynamical stability and instability of stationary states is inferred based on two pieces of evidence: (i) the negative slope of the $(I_{\mathrm{vc,sm}},V_{\mathrm{h}})$-curve (brown) after smoothing over a moving window with larger size ($\Delta w=1.5\times 10^{4}$ steps equaling =0.75=0.75s in Fig 1C), or (ii) by the presence of slow-fast oscillations (neuronal firing) in the CC run at $I_{\mathrm{h}}$ equalling the $I_{\mathrm{vc,sm}}$ (red). Criterion (i) implies instability for topological reasons. Criterion (ii) provides only circumstantial evidence. In Fig 1B (cell 5) the darker shading of the (orange-colored) CC run indicates slow dynamics as identified by the norm $|(V_{\mathrm{cc}}^{\prime}(t)\sqrt{\Delta_{t}},V_{\mathrm{cc}}^{\prime\prime}(t))\Delta_{t}|$ being less than a fixed threshold (after smoothing of $V_{\mathrm{cc}}(t)$ and $V_{\mathrm{cc}}^{\prime}(t)$, see S1 Text for detailed definition). This shading shows that the firing oscillations in Fig 1B spend most time near their voltage minimum $V_{\min}$ for the respective $I_{\mathrm{h}}(t)$. Fig 5C and 5E show an embedding of the firing oscillations of cell 5 into the $(V_{\mathrm{cc}},V_{\mathrm{cc}}^{\prime})$-plane, and a zoom near $I_{\mathrm{h}}\approx 200$ pA. The projection and zoom indicate that the firing oscillation passes slowly near the presumed fixed point, but leaves its neighborhood again. At the low-current end of the $(I_{\mathrm{vc,sm}},V_{\mathrm{h}})$-curve, the CC protocol for cell 5 shows no firing oscillations (recall that $I_{\mathrm{h}}$ is ramped up) in Fig 1B. Thus, the part of the $(I_{\mathrm{vc,sm}},V_{\mathrm{h}})$ curve between fold and presence of slow-fast oscillations is labelled “undetermined” (colored yellow) as the combination of single-ramp VC and CC protocol do not provide evidence for or against stability of this part. We do not label the “transition” from stability label “undetermined” to ”unstable” as a Hopf bifurcation as the slow parts of the firing oscillations are approximately $20$ mV below the equilibrium indicated by the VC run. So, it is unclear if (and where precisely) a change of stability occurs between $I_{\mathrm{db}}\approx 365$ pA and the fold of steady states at $I_{\mathrm{h}}\approx 3$ pA.
@@ 2 Results
The stability boundaries of the stationary states are labelled as bifurcations in Fig 1B. The change of stability near the disappearance of stable spiking states at $I_{\mathrm{db}}$ is labelled as a Hopf bifurcation. We observe that for $I_{\mathrm{h}}>I_{\mathrm{db}}$ small-amplitude oscillations are visible, which emerge from relaxation type oscillations with a slow phase near the stationary state for $I_{\mathrm{h}}<I_{\mathrm{db}}$. Fig I in S1 Text shows a zoom of Fig 1B near $I_{\mathrm{db}}$. These features are typical for a singular Hopf bifurcation [baer1986singular] as encountered also in model simulations for excitatory neurons. The fold points of the $(I_{\mathrm{vc,sm}},V_{\mathrm{h}})$-curve are saddle-node (fold) bifurcations. Fig 1C shows the stationary-state curves with their inferred dynamical stability for all $5$ cells (see S1 Text for CC run time profiles used to partially infer stability of cells 1144). The cells are vertically ordered and numbered according to depth of the S-shape, determining if they fall into the category of class-I or class-II neurons. We use the same convention as in Fig 1B for parts of the curve where we have partial evidence for instability: no bifurcation is indicated at transitions between stability labels “undetermined” and “unstable”. Fig 1C demonstrates wide variability in steady-state curve shape among cells of nominally same function.
@@
Figure 2: A: Experimental bifurcation diagram for an interneuron from the entorhinal cortex. B: Experimental bifurcation diagram for a class-II PY neuron from the same region. Protocols identical and color coding to Fig 1B.
The method described in this work is versatile and uses only standard electrophysiological procotols. Hence, it can be applied to any neuron type. The experimental bifurcation diagrams in Fig 2A illustrate this point. Panel A shows the results of the VC and CC protocols applied to an inhibitory neuron from the entorhinal cortex while panel B shows the results of same protocols applied to a class-II neuron. The steady-state curve in Fig 2A is more complicated than for PY neurons’ recordings shown in other figures. However it is compatible with a class-I excitable neuron. Since the steady-state curve in Fig 2A has multiple folds, there are some parts of the branch where we cannot determine stability safely from topology (that is, by checking that $\mathrm{d}I_{\mathrm{vc,sm}}/\mathrm{d}V_{\mathrm{h}}<0$), or from the near-by embedded trajectories of relaxation oscillations. Similar to Fig 1 we have colored the parts of the steady-state curve in Fig 2A with undetermined stability yellow, while keeping the part for which instability is suggested by existence of relaxation oscillations in red. For the range of $I_{\mathrm{h}}$ between $43$ pA and $60$ pA during the CC protocol in Fig 2A we observe a stable steady state $V_{\mathrm{cc}}$ response, with intermittent voltage spikes typical for an excitable stable steady state near a singular Hopf bifurcation subject to noise [Lindner2004]. For this reason we label the low-$I_{\mathrm{h}}$ onset of large-amplitude relaxation oscillations in Fig 2A also as a Hopf bifurcation. A potential alternative mechanism for the sudden transition of firing oscillations to steady state near $I_{\mathrm{db}}$ ($\approx 210$ pA for Fig 2A, $\approx 190$ pA for Fig 2B) is that the firing oscillations experience a saddle-node of periodic orbits. This would imply a bistability between firing and steady state for $I_{\mathrm{h}}<I_{\mathrm{db}}$, a scenario that has been observed near depolarization blocks in models of dopaminergic neurons by [dovzhenok2012exploring]. Our illustrative examples of mathematical neuron models shown in Figs 3A and 4 have such a small region of bistability (Fig 3A near $I_{\mathrm{db}}$ and Fig 4 near $I_{\mathrm{h}}\approx 135$ pA) for the parameters used in the computations (see Figs F, G and H in S1 Text for numerical bifurcation diagrams).
We observe that the steady-state curves for the CC run and the VC run slightly deviate from each other, which we attribute to a drift in cell properties. There are examples of scenarios with multiple folds in the steady-state curve in the literature, in particular with low-threshold spiking neurons [rush1994].
2.1 Analysis
@@
describing the current balance across the neuron’s membrane. The membrane potential is $V$, $I_{j}(x_{j},V)_{j=1,\ldots,N}$ are the currents across different voltage-gated ion channel types and $I_{\mathrm{ext}}$ is the external current. Each channel type $j$ has a set of associated gating variables $x_{j}$, which are possibly vectors of length $n_{j}$ if the channel gate has both activation and inactivation states, with steady-state gating functions $x_{\infty,j}(V)$ (also of size $n_{j}$) and relaxation times $\tau_{j}(V)$ (a diagonal $n_{j}\times n_{j}$ matrix). The observed dynamic effects such as oscillations (firing/spiking) and negative-slope $(I,V)$ characteristics are determined by these channel coefficients $I_{j}$, $x_{\infty,j}$ and $\tau_{j}$ that are traditionally obtained by parameter fitting from VC experiments, a difficult and ill-posed problem, as gating variables $x_{j}$ are not directly measured [de2018].
@@
The VC and CC protocols use different mechanisms for generating $I_{\mathrm{ext}}(t)$. The VC protocol is a closed loop where a voltage source regulates $I_{\mathrm{ext}}$ with high gain $g_{\mathrm{c}}$ to achieve the slowly varying hold voltage $V_{\mathrm{h}}$ at the voltage source for general model (1), measuring $I_{\mathrm{vc}}$:
$I_{\mathrm{ext}}$ $\approx I_{\mathrm{vc}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V),$ $\dot{V}_{\mathrm{h}}$ $=\varepsilon\Delta_{V}$ (2)
@@
which turns general model (1) with VC protocol (2) into a multiple-timescale dynamical system with $1+\sum n_{j}$ fast state variables $(V,x_{j})$ (where $\sum n_{j}$ is the overall number of gating variables) and one slow state variable $V_{\mathrm{h}}$, corresponding to the feedback reference signal [izhikevich2007]. The speed at which $V_{\mathrm{h}}$ varies is $\varepsilon\Delta_{V}$ with $\Delta_{V}(t)=0.183$ mV//ms, where we extract the dimensionless small factor $\varepsilon=10^{-2}$.
@@
In contrast, the CC protocol holds $I_{\mathrm{ext}}$, measuring the generated voltage $V_{\mathrm{cc}}$, thus, corresponding to an open-loop system, permitting e.g. the spiking seen in Fig 1B:
$I_{\mathrm{ext}}$ $\approx I_{\mathrm{h}},$ $V_{\mathrm{cc}}$ $\approx V,$ $\dot{I}_{\mathrm{h}}$ $=\varepsilon\Delta_{I}.$ (3)
@@
The applied hold current $I_{\mathrm{h}}$ is varied slowly at speed $\varepsilon\Delta_{I}$ with $\varepsilon=10^{-2}$ and $\Delta_{I}=0.66\ldots 0.75$ pA//ms. General model (1) with CC protocol (3) is also a slow-fast system with $1+N$ fast variables $(V,x_{j})$ and $1$ slow variable $I_{\mathrm{h}}$. The gain $g_{\mathrm{c}}$ in VC protocol (2) is limited by the imperfect conductance across the non-zero spatial extent of the membrane. Even though the conductance-based model (1) is for the potential $V$ across the entire membrane and only $V_{\mathrm{h}}$ at the clamp is measured, we approximate $V_{\mathrm{h}},V_{\mathrm{cc}}\approx V$ for the membrane potential, and $I_{\mathrm{h}},I_{\mathrm{vc}}\approx I_{\mathrm{ext}}$ for the external current in (1).
@@
Following a classical multiple-timescale approach, we consider the $\varepsilon=0$ limit of general model (1) with VC protocol (2), which corresponds to its $(1+\sum n_{j})$-dimensional fast subsystem (1), with $I_{\mathrm{ext}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$, where $V_{\mathrm{h}}$ is now treated as a parameter. For fixed $V_{\mathrm{h}}$ and voltages in the range $80\ldots 30$ mV of interest, model (1) with VC protocol (2) has only stable steady states (no limit cycles, that is, no neuronal spikes); see S1 Text for details. For a fixed hold voltage $V_{\mathrm{h}}$, the steady states of system (1), (2) with $I_{\mathrm{ext}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$ satisfy the algebraic equations for $(V,x_{j})_{j=1,\ldots,N}$
$\textstyle{\sum_{j}}I_{j}(x_{j},V)$ $=g_{\mathrm{c}}(V_{\mathrm{h}}-V),$ $x_{j}$ $=x_{\infty,j}(V),$ (4)
@@
where $I_{\mathrm{eq}}(V)=\sum_{j}I_{j}(x_{\infty,j},V)$ is the equilibrium current for fixed membrane potential $V$. The solutions of algebraic system (4) form a (1D) steady-state curve for model (1) with VC protocol (2), which is normally hyperbolic (transversally attracting) for $\varepsilon=0$. For $\varepsilon\neq 0$ the increase of $V_{\mathrm{h}}$ with speed $\varepsilon\Delta_{V}$ introduces a slow variation of all states $(V,x_{j})$. Hence, $V$ and the feedback current $I_{\mathrm{vc}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$ (as measured), are not at their steady-state values given by (4), but they are changing dynamically. This results in a difference between the measured curve $(I_{\mathrm{vc}},V_{\mathrm{h}})$ in Fig 1B and the $(I,V)$-values of the desired steady-state IVI-V curve given by (4).